OpenStudy (anonymous):

complete the square -x^2+7x-13

OpenStudy (solomonzelman):

\[-x^2+7x-13=0\]

OpenStudy (solomonzelman):

1) multiply everything times -1. 2) (since 7=3.5 * 2 and 3.5^2=12.25 ) subtract .75 from both sides If you haven't learned about imaginary numbers, than say no solution.

OpenStudy (anonymous):

Divided both sides by negative 1 to get X^2-7x+13=0 X^2-7x+42.25+13=42.25 X^2-7x+55.25=42.25 ?????????

OpenStudy (solomonzelman):

No, lightning \[-x^{2}+7x-13=0\]\[x^{2}-7x+13=0\]\[x^{2}-7x+13-0.75=0-0.75\]\[x^{2}-7x+12.25=-.75\]\[(x-3.5)^{2}=-.75\] the solutions will be imaginary.

OpenStudy (solomonzelman):

@ashleylk5 , which course is this? Algebra 1? In your class, have you learned about imaginary numbers, or about i that is equal to √-1 ?

OpenStudy (anonymous):

Thanks you guys ^^

OpenStudy (solomonzelman):

You welcome:) Do you need more help ? reply if you do... Also, for a future reference: you can find any section by: 1) clicking "more subjects" ( top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying. you can fan someone by: 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" ) you can mention someone in your question (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @ashleylk5 (they would get a notification that they got mentioned just like you did) you can message someone by: 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. If you have any questions you can message me or anyone else on this website. We will be happy to respond !